OpenStudy (anonymous):

If height is increasing at a rate of 5m/s while radius is constant, Find the rate of increase of volume of cylinder

OpenStudy (anonymous):

@nincompoop I just want how to proceed

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

Start by giving a volume to your cylinder ok? Make it an easy #. Remember the units will be cubed, in this case m^3

OpenStudy (anonymous):

V=pir^2h

OpenStudy (anonymous):

make it a little fast , i will understand don't worry

OpenStudy (anonymous):

Icky problem, I need a sec here. lol

OpenStudy (anonymous):

Say your height is 1m and your radius is 1m then your V = pi(1^2)(1) or V = pi

OpenStudy (anonymous):

After 1 second your h is now 6m, and your r is still 1m

OpenStudy (anonymous):

Calculate your V. When you started your volume was at t=0s or V(t1) With the last Calculation your V(t2) or 1s is ?

OpenStudy (anonymous):

This is calculus problem actually

OpenStudy (anonymous):

11s at (t3) etc. Use the rate of change from initial time to final time from t3 to t2 or t1 to t0 to calculate your slope on a graph or represented rate of change.

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

Sorry about that if I led you astray...

OpenStudy (anonymous):

Take my medal away please. I did not earn that one haha